OpenStudy (anonymous):

[(7x+5)+4y][(7x+5)-4y]

OpenStudy (anonymous):

\[[a-b][a+b]=a^2-b^2\]

OpenStudy (anonymous):

(a+b) (a-b)=a2-b2 then put 7x+5 in place of a and 4y in place of b.... then calculate and get the answer

OpenStudy (anonymous):

That would only remove the grouping signs though, right?

mathslover (mathslover):

yes but you will "have" to simplify it further also

OpenStudy (anonymous):

So it is written like 7x+5+4y=7x+5^2-4y^2

mathslover (mathslover):

\[\large{(2x+5)^2-(4y)^2}\]

mathslover (mathslover):

You have to simplify it further

mathslover (mathslover):

\[\large{(7x+5)^2-(4y)^2}\] \[\large{(a+b)^2=a^2+2ab+b^2}\] \[\large{(7x+5)^2=? }\]

OpenStudy (anonymous):

 this is the answer but u have to remember the formula that (a+b) (a-b)=a square - b square. this formula will help u in many other poblems